Mrs. Diaz wishes to save at least $1,500 in 12 months. If she saves $300 during the first 4 months, what is the least possible average amount that she must save in each of th remaining 8 months?
Answer by rfer(16322)

300*4=1200
1500-1200=300/8=$37.50 per mo